OpenStudy (anonymous):

how to find the equation of the line where two planes meet? (row operations help please!)

OpenStudy (anonymous):

P1: 2x –y +4z = 14 P2: 3x +5y + 6z = 21

OpenStudy (anonymous):

ive reduced it down to 1 | -0.5 | 2 : 7 0 | 1 | 0 : 0

OpenStudy (anonymous):

therefore y=0

OpenStudy (anonymous):

but im not sure what to sub next

OpenStudy (anonymous):

i have to show the equaiton of the line is x=3-2t y=0 z=2+t

ganeshie8 (ganeshie8):

taking the cross product of normal vectors of given planes gives you the direction vector of intersection right ?

OpenStudy (anonymous):

yeah i could do that.....

OpenStudy (anonymous):

a cross b = [-26, 0, 13]

OpenStudy (anonymous):

and i have two points on the line A(3, 0, 2) B(1, 0, 3)

OpenStudy (anonymous):

and a cross b = [-2, 0, 1]

OpenStudy (anonymous):

voila - i just have to use point a

OpenStudy (anonymous):

thanks @ganeshie8

ganeshie8 (ganeshie8):

yes :) direction vector and a point is sufficient to fully define a line

ganeshie8 (ganeshie8):

row reduction should also work

OpenStudy (anonymous):

....it didn't seem to...

ganeshie8 (ganeshie8):

just express the pivot variables in terms of free variables

OpenStudy (anonymous):

okay..... i got y=0 then x+2z = 7

OpenStudy (anonymous):

so let z=t

OpenStudy (anonymous):

then x=7-2t ... which isn't right?

ganeshie8 (ganeshie8):

2x –y +4z = 14 3x +5y + 6z = 21 -------------------------- say the direction vector of required plane = <p,q,r> since the required line is perpendicular to both the planes, the dot product of direction vector with the normal vectors of planes is 0 : 2p-q++4r = 0 3p+5q+6r = 0

ganeshie8 (ganeshie8):

you need to solve above homogeneous system right ?

OpenStudy (anonymous):

ooh i didn't think of it that way....

ganeshie8 (ganeshie8):

the matrix equation would be : \[ \left[\begin{array}\\ 2&-1&4 \\ 3&5&6 \\ \end{array}\right] \left[\begin{array}\\ p\\ q\\ r \end{array}\right] = \left[\begin{array}\\ 0\\ 0\\ \end{array}\right] \]

ganeshie8 (ganeshie8):

you can row reduce and the infinitely many solutions give you the direction vector of intersecting line

ganeshie8 (ganeshie8):

thats the idea right ?

ganeshie8 (ganeshie8):

the matrix equation would be : \[ \left[\begin{array}\\ 2&-1&4 \\ 3&5&6 \\ \end{array}\right] \\~\\ \left[\begin{array}\\ 2&-1&4 \\ 0&13/2&0 \\ \end{array}\right] \\~\\ \left[\begin{array}\\ 2&-1&4 \\ 0&1&0 \\ \end{array}\right]~\\~\\ \left[\begin{array}\\ 2&0&4 \\ 0&1&0 \\ \end{array}\right]~\\~\\ \left[\begin{array}\\ 1&0&2 \\ 0&1&0 \\ \end{array}\right]~\\~\\ \]

ganeshie8 (ganeshie8):

p = -2 q = 0 r = 1 as expected, eh ?

ganeshie8 (ganeshie8):

there was a serious typo in my earlier reply, corrected below : 2x –y +4z = 14 3x +5y + 6z = 21 -------------------------- say the direction vector of require \(\color{red}{\text{line}}\) = <p,q,r> since the required line is perpendicular to both the planes, the dot product of direction vector with the normal vectors of planes is 0 : 2p-q++4r = 0 3p+5q+6r = 0